OpenStudy (anonymous):

What is the value of the expression when a = 4 and b = 2? 2a^3+2b^2

Vocaloid (vocaloid):

a = 4 b = 2 means you can replace "a" with 4 and "b" with 2

Vocaloid (vocaloid):

so if a = 4, what is 2a^3?

OpenStudy (anonymous):

24^3

Vocaloid (vocaloid):

good, so 2a^3 = 2(4^3) which equals...?

OpenStudy (anonymous):

B.80?

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

what is 4^3?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

nope, remember that ^ means an exponent

Vocaloid (vocaloid):

4^3 means we multiply 4 by itself 3 times so 4^3 = 4*4*4 = ?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

good, now what is 2(64)?

OpenStudy (anonymous):

128

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

good. now, if b = 2, what is 2(b^2)?

OpenStudy (anonymous):

2*2=4

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

So my answer is C.120?

Vocaloid (vocaloid):

nope

OpenStudy (anonymous):

Because, 116+4=120

Vocaloid (vocaloid):

2*b^2 = 2*b*b = ?

OpenStudy (anonymous):

2*b^2=8

OpenStudy (anonymous):

right?

Vocaloid (vocaloid):

good. now for the first part we got 128 and for the second part we got 8 now add them together

OpenStudy (anonymous):

Ahhhhh Makes sense. 136. The answer is D